# Optimal Data Splitting for Holdout Cross-Validation in Large Covariance Matrix Estimation

### v1

@@ 2.2.2 Cross-validation estimators
In this section, we define the two main CV methods: the holdout method and the $k$-fold CV. Additionally, we introduce two variations of these methods, one with and one without rotational invariance. To explain these methods, we need first to define the index partitions.
@@
Let $t$ be the total number of observations, and let $t_{\textrm{out}}$ be an integer between $1$ and $t-1$. Define $t_{\textrm{in}}=t-t_{\textrm{out}}$. We partition the set of indices =$\mathcal{I}=\{1,2,\dots,t\}$ into $2$ disjoint subsets $\mathcal{I}_{\textrm{out}}$ and $\mathcal{I}_{\textrm{in}}$ of size $t_{\textrm{out}}$ and $t_{\textrm{in}}$ respectively, such that
@@
Let $t$ be the total number of observations, and let $t_{\textrm{out}}$ be an integer between $1$ and $t-1$ such that $t_{\textrm{out}}$ divides $t$. Define $k=t/t_{\textrm{out}}$ and $t_{\textrm{in}}=t-t_{\textrm{out}}$. We partition the set of indices =$\mathcal{I}=\{1,2,\dots,t\}$ into $k$ disjoint subsets $\{\mathcal{I}_{\textrm{out},l}\}_{l=1}^{k}$, each of size $t_{\textrm{out}}$, such that
@@
Let us consider $\bm{X}=\{\bm{x}_{1},...,\bm{x}_{t}\}$ an i.i.d. set of centered observations of covariance $\bm{\Sigma}$. Let $t_{\textrm{out}}$ an integer between $1$ and $t-1$, and $\{\mathcal{I}_{\textrm{in}},\mathcal{I}_{\textrm{out}}\}$ an index partition (definition  (2.11)) defined from $t_{\textrm{out}}$. Let us call $\bm{E}_{\textrm{in}}$ the sample covariance over the subset of observations $\bm{X}_{\textrm{in}}=\{\bm{x}_{i}|\,i\in\mathcal{I}_{\textrm{in}}\}$ (train set) and $\bm{E}_{\textrm{out}}$ the sample covariance on the remaining observations $\bm{X}_{\textrm{out}}=\{\bm{x}_{i}|\,i\in\mathcal{I}_{\textrm{out}}\}$ (test set). Let us call $\bm{V}_{\textrm{in}}$ the eigenvectors of $\bm{E}_{\textrm{in}}$ ranked by increasing eigenvalues. Then the holdout estimator is
@@
Let us consider $\bm{X}=\{\bm{x}_{1},...,\bm{x}_{t}\}$ an i.i.d. set of centered observations of population covariance $\bm{\Sigma}$ and sample covariance $\bm{E}$. Let $t_{\textrm{out}}$ an integer between $1$ and $t-1$,and $\{\mathcal{I}_{\textrm{in}},\mathcal{I}_{\textrm{out}}\}$ an index partition (definition  (2.11)) defined from $t_{\textrm{out}}$. Let us call $\bm{E}_{\textrm{in}}$ the sample covariance over the subset of observations $\bm{X}_{\textrm{in}}=\{\bm{x}_{i}|\,i\in\mathcal{I}_{\textrm{in}}\}$ (train set) and $\bm{E}_{\textrm{out}}$ the sample covariance on the remaining observations $\bm{X}_{\textrm{out}}=\{\bm{x}_{i}|\,i\in\mathcal{I}_{\textrm{out}}\}$ (test set). Let us call $\bm{V}_{\textrm{in}}$ and $\bm{V}$ the eigenvectors of $\bm{E}_{\textrm{in}}$ and $\bm{E}$ respectively ranked according to their eigenvalues. Then the holdout eigenvalues are
@@
Let us consider $\bm{X}=\{\bm{x}_{1},...,\bm{x}_{t}\}$ an i.i.d. set of centered observations of covariance $\bm{\Sigma}$. Let $t_{\textrm{out}}$ be an integer between $1$ and $t-1$ such that $t_{\textrm{out}}$ divides $t$, and let $\{\mathcal{I}_{\textrm{out},l},\mathcal{I}_{\textrm{in},l}\}$ be a $k$-fold index partition (definition (2.12)). For each $1\leq l\leq k$, let $\bm{E}_{\textrm{in},l}$ be the sample covariance over the observations $\bm{X}_{\textrm{in},l}=\{\bm{x}_{i}|i\in\mathcal{I}_{\textrm{in},l}\}$ (train set) and $\bm{E}_{\textrm{out},l}$ the sample covariance over the observations $\bm{X}_{\textrm{out},l}=\{\bm{x}_{i}|i\in\mathcal{I}_{\textrm{out},l}\}$ (test set). Let $\bm{V}_{\textrm{in},l}$ the eigenvectors of $\bm{E}_{\textrm{in},l}$ ranked by increasing eigenvalues. Then the $k$-fold CV estimator is
@@
Let us consider $\bm{X}=\{\bm{x}_{1},...,\bm{x}_{t}\}$ an i.i.d. set of centered observations of population covariance $\bm{\Sigma}$ and sample covariance $\bm{E}$. Let $t_{\textrm{out}}$ be an integer between $1$ and $t-1$ such that $t_{\textrm{out}}$ divides $t$, and let $\{\mathcal{I}_{\textrm{in},l},\mathcal{I}_{\textrm{out},l}\}_{l=1}^{k}$ be a $k$-fold index partition (definition (2.12)). For each $1\leq l\leq k$, let $\bm{E}_{\textrm{in},l}$ be the sample covariance over the observations $\bm{X}_{\textrm{in},l}=\{\bm{x}_{i}|i\in\mathcal{I}_{\textrm{in},l}\}$ (train set) and $\bm{E}_{\textrm{out},l}$ the sample covariance over the observations $\bm{X}_{\textrm{out},l}=\{\bm{x}_{i}|i\in\mathcal{I}_{\textrm{out},l}\}$ (test set). Let $\bm{V}_{\textrm{in},l}$ and $\bm{V}$ be eigenvectors of $\bm{E}_{\textrm{in},l}$ and $\bm{E}$ respectively ranked according to their eigenvalues. Then the $k$-fold CV eigenvalues are
@@ 3 Derivation of the Estimation Error for the Holdout
Before carrying on the derivation of the error, it is worth having a look at the shapes of the Frobenius error as a function of the as a function of the train-test ratio factor $k=t/t_{\textrm{out}}$. In figure 1, we show that either CV and Holdout methods exhibit a minimum error point that is not at an extreme value of $k$. Although the exact locations of the minimum can be different among the two methods, numerical simulations seem to indicate a similar scale factor. This discrepancy might arise from the distinct meaning of $k$ in each case: in the CV method, $k$ represents the number of splits, while in the Holdout method, it reflects only the imbalance between the sizes of the training and testing sets.
@@ 3.1 General case
The proof involves two steps. The first one is to apply Wick’s theorem to link the moments of the sample covariances to the population covariance matrix using the Gaussian assumption [42]. The second step is to recognize the oracle estimator eigenvalues for which we have a simple expression in the inverse Wishart case using the Ledoit-Péché’s formula [16].
@@
Let us consider $\bm{\Sigma}$ a positive-defined matrix in $\mathbb{R}^{n\times n}$ and $(\bm{x}_{\textrm{i}})_{1\leq i\leq t}$ i.i.d. random vectors of $\mathbb{R}^{n}$ such that $\bm{x}_{i}\sim\mathcal{N}(0,\bm{\Sigma})$. Let $\bm{\Xi}^{\textrm{H}}$ the holdout estimator associated with the index partition $\{\mathcal{I}_{\textrm{in}},\mathcal{I}_{\textrm{out}}\}$ defined from $t_{\textrm{out}}$ integer between $1$ and $t-1$, with $\bm{V}_{\textrm{in}}$ the eigenvector of the sample covariance $\bm{E}_{\textrm{in}}$ on the train partition $\mathcal{I}_{\textrm{in}}$. Then the expected Frobenius error between the holdout estimator and the population covariance $\bm{\Sigma}$ is
@@
Let us consider $\{\bm{x}_{1},\bm{x}_{2},..,\bm{x}_{r}\}$ i.i.d. centered Gaussian vectors of $\mathbb{R}^{n}$, then [42]
@@
From this expression, we recognize the oracle eigenvalues $\bm{\Lambda}^{O}$ of equation (20). In the high dimension limit, we can use the Ledoit Péché formula (equation 21) for $\bm{\Lambda}^{O}$. The formula is exact in the high dimension limit but remains a very good approximation when the dimensions, $n$ and $t$ are finite but large [17]. The formula in equation (40) is the general holdout error for Gaussian multiplicative noise and establishes a quantitative link with the oracle estimator eigenvalues and it is not restricted to the inverse Wishart case. One can compute the holdout error for any population covariance matrix as long as one can apply the Ledoit-Péché formula written in equation (21). Therefore when $\bm{\Sigma}$ has a spectral density which converges a.s. at every point of continuity to a non-random limit when $n$ tends to $\infty$ and so that the support of the limiting distribution is included in a compact interval [11]. Moreover, this equation can be solved numerically for any population covariance verifying these regularity conditions [17, 18].
@@ 3.2 White inverse Wishart case
Let us consider $\bm{\Sigma}\sim\mathcal{W}^{-1}_{np}$ a white inverse Wishart of parameters $(n,p)$, with $p$ negligible relative to $n$, and $(\bm{x}_{\textrm{i}})_{1\leq i\leq t}$ i.i.d. random vectors of $\mathbb{R}^{n}$ such that $\bm{x}_{i}\sim\mathcal{N}(0,\bm{\Sigma})$. Let $\bm{\Xi}^{\textrm{H}}$ the holdout estimator associated with the index partition $\{\mathcal{I}_{\textrm{in}},\mathcal{I}_{\textrm{out}}\}$ defined from $t_{\textrm{out}}$ integer between $1$ and $t-1$. The expected Frobenius error between the holdout estimator and the population covariance $\bm{\Sigma}$ is
@@
To prove it, we first note that
Proposition 3.3.
When the population matrix is a white inverse Wishart, the oracle estimator recovers the optimal linear shrinkage in the Bayesian sense, in the high dimension limit when $p$ is negligible compared to $n$ and is equal to [16]
@@
Since the data are Gaussian, one can apply lemma (2.2) and proposition (2.2) if $p$ is negligible relative to $n$, then
@@
and being the population matrix a white inverse Wishart, from the definition (2.5) and lemma (2.1), we have
@@
In the left panel of figure 2, we compared the theoretical error formula equation (45) with extensive Monte Carlo estimations. The theoretical error shows a good agreement with the data when $p$ is negligible compared to $n$. However, when $p$ is larger compared to $n$, the error formula shows a systematic underestimation of the error. This behavior appears clear since only the points associated with $p/n>10^{-2}$ are biased. The reason is probably that both in the optimal linear shrinkage and in our derivation, equation 12, requires $p$ significantly smaller than $n$.
In the right panel of figure 2, we show the theoretical and Monte Carlo error as a function of $k$ for a specific set of $\{n,p,q\}$. Interestingly, even though for our choice of parameters, a small bias is still observable, the position of the minimum seems not affected. This motivates us to use equation (45) to derive an analytical expression for the optimal $k$.
Figure 2: The left panel shows the comparison of the theoretical error of equation (45) with a Monte Carlo estimations for a random uniform selection of the parameters: $n\in[100,1000]$, $p\in[0.1,9]$, $q\in[0.1,0.9]$ and $k$ among the divisors of $t$. For stability, every combination of parameters is repeated $100$ times, and the estimated error is averaged. The right panel shows the theoretical Frobenius error and the Monte Carlo estimation as a function of $k$ for $n=750$, $t=1000$, and $p=0.06$ averaged $1000$ times.

### v2

@@ 2.2.2 Cross-validation estimators
In this section, we define the two main CV methods: the holdout method and the $k$-fold CV. Additionally, we introduce two variations of these methods, one with and one without rotational invariance (Abadir et al., 2014; Lam, 2016). To explain these methods, we need first to define the index partitions.
@@
Let $t$ be the total number of observations, and let $t_{\textrm{out}}$ be an integer between $1$ and t1t-1. Define $t_{\textrm{in}}=t-t_{\textrm{out}}$. We partition the set of indices $\mathcal{I}=\{1,2,\dots,t\}$ into $2$ disjoint subsets $\mathcal{I}_{\textrm{out}}$ and $\mathcal{I}_{\textrm{in}}$ of size $t_{\textrm{out}}$ and $t_{\textrm{in}}$ respectively, such that
@@
Let $t$ be the total number of observations, and let $t_{\textrm{out}}$ be an integer between $1$ and t1t-1 such that $t_{\textrm{out}}$ divides $t$. Define $k=t/t_{\textrm{out}}$ and $t_{\textrm{in}}=t-t_{\textrm{out}}$. For the k-fold, we partition the set of indices $\mathcal{I}=\{1,2,\dots,t\}$ into $k$ disjoint subsets $\{\mathcal{I}_{\textrm{out},l}\}_{l=1}^{k}$, each of size $t_{\textrm{out}}$, such that
@@
Let us consider $\bm{X}=\{\bm{x}_{1},...,\bm{x}_{t}\}$ an i.i.d. set of centered observations of covariance $\bm{\Sigma}$. Let $t_{\textrm{out}}$ an integer between $1$ and t1t-1, and $\{\mathcal{I}_{\textrm{in}},\mathcal{I}_{\textrm{out}}\}$ an index partition (definition  (2.7)) defined from $t_{\textrm{out}}$. Let us call $\bm{E}_{\textrm{in}}$ the sample covariance over the subset of observations $\bm{X}_{\textrm{in}}=\{\bm{x}_{i}|\,i\in\mathcal{I}_{\textrm{in}}\}$ (train set) and $\bm{E}_{\textrm{out}}$ the sample covariance on the remaining observations $\bm{X}_{\textrm{out}}=\{\bm{x}_{i}|\,i\in\mathcal{I}_{\textrm{out}}\}$ (test set). Let us call $\bm{V}_{\textrm{in}}$ the eigenvectors of $\bm{E}_{\textrm{in}}$ ranked by increasing eigenvalues. Then the holdout estimator is (Lam, 2016)
@@
Let us consider $\bm{X}=\{\bm{x}_{1},...,\bm{x}_{t}\}$ an i.i.d. set of centered observations of population covariance $\bm{\Sigma}$ and sample covariance $\bm{E}$. Let $t_{\textrm{out}}$ an integer between $1$ and t1t-1,and $\{\mathcal{I}_{\textrm{in}},\mathcal{I}_{\textrm{out}}\}$ an index partition (definition  (2.7)) defined from $t_{\textrm{out}}$. Let us call $\bm{E}_{\textrm{in}}$ the sample covariance over the subset of observations $\bm{X}_{\textrm{in}}=\{\bm{x}_{i}|\,i\in\mathcal{I}_{\textrm{in}}\}$ (train set) and $\bm{E}_{\textrm{out}}$ the sample covariance on the remaining observations $\bm{X}_{\textrm{out}}=\{\bm{x}_{i}|\,i\in\mathcal{I}_{\textrm{out}}\}$ (test set). Let us call $\bm{V}_{\textrm{in}}$ and $\bm{V}$ the eigenvectors of $\bm{E}_{\textrm{in}}$ and $\bm{E}$ respectively ranked according to their eigenvalues. Then the holdout eigenvalues are (Abadir et al., 2014)
@@
Let us consider $\bm{X}=\{\bm{x}_{1},...,\bm{x}_{t}\}$ an i.i.d. set of centered observations of covariance $\bm{\Sigma}$. Let $t_{\textrm{out}}$ be an integer between $1$ and t1t-1 such that $t_{\textrm{out}}$ divides $t$, and let $\{\mathcal{I}_{\textrm{out},l},\mathcal{I}_{\textrm{in},l}\}$ be a $k$-fold index partition (definition (2.8)). For each $1\leq l\leq k$, let $\bm{E}_{\textrm{in},l}$ be the sample covariance over the observations $\bm{X}_{\textrm{in},l}=\{\bm{x}_{i}|i\in\mathcal{I}_{\textrm{in},l}\}$ (train set) and $\bm{E}_{\textrm{out},l}$ the sample covariance over the observations $\bm{X}_{\textrm{out},l}=\{\bm{x}_{i}|i\in\mathcal{I}_{\textrm{out},l}\}$ (test set). Let $\bm{V}_{\textrm{in},l}$ the eigenvectors of $\bm{E}_{\textrm{in},l}$ ranked by increasing eigenvalues. Then the $k$-fold CV estimator is (Lam, 2016)
@@
Let us consider $\bm{X}=\{\bm{x}_{1},...,\bm{x}_{t}\}$ an i.i.d. set of centered observations of population covariance $\bm{\Sigma}$ and sample covariance $\bm{E}$. Let $t_{\textrm{out}}$ be an integer between $1$ and t1t-1 such that $t_{\textrm{out}}$ divides $t$, and let $\{\mathcal{I}_{\textrm{in},l},\mathcal{I}_{\textrm{out},l}\}_{l=1}^{k}$ be a $k$-fold index partition (definition (2.8)). For each $1\leq l\leq k$, let $\bm{E}_{\textrm{in},l}$ be the sample covariance over the observations $\bm{X}_{\textrm{in},l}=\{\bm{x}_{i}|i\in\mathcal{I}_{\textrm{in},l}\}$ (train set) and $\bm{E}_{\textrm{out},l}$ the sample covariance over the observations $\bm{X}_{\textrm{out},l}=\{\bm{x}_{i}|i\in\mathcal{I}_{\textrm{out},l}\}$ (test set). Let $\bm{V}_{\textrm{in},l}$ and $\bm{V}$ be eigenvectors of $\bm{E}_{\textrm{in},l}$ and $\bm{E}$ respectively ranked according to their eigenvalues. Then the $k$-fold CV eigenvalues are (Abadir et al., 2014)
@@
Before considering the computation of the expected error of estimation of the holdout estimator, let us first state the assumptions of the convergence result of Lam (2016, 2015) for the holdout error towards the NLS error.
Assumption 2.1.
The observations can be written as $\bm{X}=\sqrt{\bm{\Sigma_{n}}}\bm{Y}$ where $\bm{Y}$ is a $n\times t$ matrix of real or complex random variables with zero mean and unit variance such that $\mathbb{E}\left(|Y_{ij}|^{l}\right)\leq b<\infty$ for some constant $b$ and for $2<l\leq 20$.
The population covariance matrix $\bm{\Sigma}_{n}$ is non-random, $\|\bm{\Sigma}\|_{L^{2}}=\mathcal{O}(1)$ and $\bm{\Sigma}\neq\sigma^{2}\bm{\mathds{1}}$
$\frac{n}{t}\to q>0$ as $n\to\infty$
Let us consider $(\tau_{1}\dots\tau_{n})$ the eigenvalues of $\bm{\Sigma}_{n}$, the empirical spectral distribution of $\bm{\Sigma}_{n}$
$H_{n}(\tau)=\frac{1}{n}\sum_{j=1}^{n}\mathds{1}_{[\tau_{j},+\infty[}(\tau)$ (32)
converges to a non-random limit $H(\tau)$ at every point of continuity of $H$. Moreover, H defines a probability distribution function whose support $Supp(H)$ is included in a compact interval $h_{1},h_{2}]$ where $0<h_{1}\leq h_{2}<+\infty$.
Let us now state Lam (2016)’s convergence result in the asymptotic regime
Theorem 2.2.
Under the assumptions listed in 2.1, the holdout estimator error converges towards the error of the NLS, with respect to the Frobenius error and Stein’s loss provided that the split satisfies $\frac{t_{\textrm{in}}}{t}\xrightarrow[t\to\infty]{}1$, $t_{\textrm{out}}\xrightarrow[t\to\infty$$\infty$ and
$\sum_{t\geq 1}\frac{n}{t_{\textrm{out}}^{5}}<\infty.$ (33)
For example, $k=\alpha\sqrt{n}$ would imply $t_{\textrm{out}}$ proportional to the square root of $n$ and would satisfy the condition on the split written in eq. (33) taking into account that $n=qt$ in the high-dimensional regime.
The last two assumptions are included in the assumptions of the Ledoit and Péché’s formula (Ledoit and Péché, 2011), moreover our setting of Gaussian data with an inverse Wishart population covariance matrix satisfies all of Lam’s assumptions except that we did not exclude the case where the population covariance matrix is proportional to the identity. Therefore, we expect to observe the asymptotic convergence of the holdout Frobenius error towards the one of the NLS.
@@ 3 Derivation of the Estimation Error for the Holdout
Before carrying on the derivation of the error, it is worth having a look at the shapes of the Frobenius error as a function of the as a function of the train-test ratio factor $k=t/t_{\textrm{out}}$. In Fig. 1, we show that either CV and Holdout methods exhibit a minimum error point that is not at an extreme value of $k$. Although the exact locations of the minimum can be different among the two methods, numerical simulations seem to indicate a similar scale factor. This discrepancy might arise from the distinct meaning of $k$ in each case: in the CV method, $k$ represents the number of splits, while in the Holdout method, it reflects only the imbalance between the sizes of the training and testing sets.
@@ 3.1 General case
The proof involves two steps. The first one is to apply Wick’s theorem to link the moments of the sample covariances to the population covariance matrix using the Gaussian assumption (Wick, 1950). The second step is to recognize the oracle estimator eigenvalues for which we have a simple expression in the inverse Wishart case using the Ledoit-Péché’s formula (Potters and Bouchaud, 2020a).
@@
Let us consider $\bm{\Sigma}$ a positive-defined matrix in $\mathbb{R}^{n\times n}$ and $(\bm{x}_{\textrm{i}})_{1\leq i\leq t}$ i.i.d. random vectors of $\mathbb{R}^{n}$ such that $\bm{x}_{i}\sim\mathcal{N}(0,\bm{\Sigma})$. Let $\bm{\Xi}^{\textrm{H}}$ the holdout estimator associated with the index partition $\{\mathcal{I}_{\textrm{in}},\mathcal{I}_{\textrm{out}}\}$ defined from $t_{\textrm{out}}$ integer between $1$ and t1t-1, with $\bm{V}_{\textrm{in}}$ the eigenvector of the sample covariance $\bm{E}_{\textrm{in}}$ on the train partition $\mathcal{I}_{\textrm{in}}$. Then the expected Frobenius error between the holdout estimator and the population covariance $\bm{\Sigma}$ is
@@
Let us consider $\{\bm{x}_{1},\bm{x}_{2},..,\bm{x}_{r}\}$ i.i.d. centered Gaussian vectors of $\mathbb{R}^{n}$, then (Wick, 1950)
@@
From this expression, we recognize the oracle eigenvalues $\bm{\Lambda}^{O}$ of equation (20). In the high dimension limit, we can use the Ledoit Péché formula (equation 21) for $\bm{\Lambda}^{O}$. The formula is exact in the high dimension limit but remains a very good approximation when the dimensions, $n$ and $t$ are finite but large (Ledoit and Wolf, 2012). The formula in equation (42) is the general holdout error for Gaussian multiplicative noise and establishes a quantitative link with the oracle estimator eigenvalues and it is not restricted to the inverse Wishart case. One can compute the holdout error for any population covariance matrix as long as one can apply the Ledoit-Péché formula written in equation (21). Therefore when $\bm{\Sigma}$ has a spectral density which converges a.s. at every point of continuity to a non-random limit when $n$ tends to $\infty$ and so that the support of the limiting distribution is included in a compact interval (Ledoit and Péché, 2011). Moreover, this equation can be solved numerically for any population covariance verifying these regularity conditions (Ledoit and Wolf, 2012, 2017).
@@ 3.2 White inverse Wishart case
Let us consider $\bm{\Sigma}\sim\mathcal{W}^{-1}_{np}$ a white inverse Wishart of parameters $(n,p)$, with $p$ negligible relative to $n$, and $(\bm{x}_{\textrm{i}})_{1\leq i\leq t}$ i.i.d. random vectors of $\mathbb{R}^{n}$ such that $\bm{x}_{i}\sim\mathcal{N}(0,\bm{\Sigma})$. Let $\bm{\Xi}^{\textrm{H}}$ the holdout estimator associated with the index partition $\{\mathcal{I}_{\textrm{in}},\mathcal{I}_{\textrm{out}}\}$ defined from $t_{\textrm{out}}$ integer between $1$ and t1t-1. The expected Frobenius error between the holdout estimator and the population covariance $\bm{\Sigma}$ is
@@
To prove it, we first note that when the population matrix is a white inverse Wishart and the data are Gaussian, the oracle estimator recovers the optimal linear shrinkage in the Bayesian sense, in the high dimension limit when $p$ is negligible compared to $n$ and is equal to (Potters and Bouchaud, 2020a)
@@
Since the data are Gaussian, one can combine equations (10) and  (14) if $p$ is negligible relative to $n$, then
@@
and being the population matrix a white inverse Wishart, from the definition (2.2) and Eq.(6), we have
@@
In the left panel of Fig. 2, we compared the theoretical error formula Eq. (47) with extensive Monte Carlo estimations. The theoretical error shows a good agreement with the data when $p$ is negligible compared to $n$. However, when $p$ is larger compared to $n$, the error formula shows a systematic underestimation of the error. This behavior appears clear since only the points associated with $p/n>10^{-2}$ are biased. The reason is probably that both in the optimal linear shrinkage and in our derivation, equation 12, requires $p$ significantly smaller than $n$.
In the right panel of Fig. 2, we show the theoretical and Monte Carlo error as a function of $k$ for a specific set of $\{n,p,q\}$. Interestingly, even though for our choice of parameters, a small bias is still observable, the position of the minimum seems not affected. This motivates us to use Eq. (47) to derive an analytical expression for the optimal $k$.
Figure 2: The left panel shows the comparison of the theoretical error of Eq. (47) with a Monte Carlo estimations for a random uniform selection of the parameters: $n\in[100,1000]$, $p\in[0.1,9]$, $q\in[0.1,0.9]$ and $k$ among the divisors of $t$. For stability, every combination of parameters is repeated $100$ times, and the estimated error is averaged. The right panel shows the theoretical Frobenius error and the Monte Carlo estimation as a function of $k$ for $n=750$, $t=1000$, and $p=0.06$ averaged $1000$ times.